Lasser Company plans to produce 16,000 units next period at a denominator activity of 32,000 direct labor-hours. The direct labor wage rate is $12.00 per hour. The company's standards allow 2 yards of direct materials for each unit of product; the material costs $7.80 per yard. The company's budget includes variable manufacturing overhead cost of $2.30 per direct labor-hour and fixed manufacturing overhead of $156,800 per period.

Required:

1. Using 32,000 direct labor-hours as the denominator activity, compute the predetermined overhead rate and break it down into variable and fixed elements.

2. Complete the standard cost card below for one unit of product.
Answer #1

1.  Predetermined overhead rate = 156800 / 32000 = $4.9 per direct labor-hour

variable manufacturing overhead cost = $2.30 per direct labor-hour

Fixed manufacturing overhead cost = $4.90 per direct labor-hour

2. STANDARD COST = 7.80*2 + 12*2 + (4.9+2.30)*2 = $54 per unit